A 2m3 volume of gas at 27.0°C is (a) heated to 327°C then (b) cooled to -123°C, both


while the pressure is constant. What are the volumes of the gas in (a) and (b)?

1
Expert's answer
2021-12-29T12:35:06-0500

According to the combined gas law:


"\frac{P_iV_i}{T_i}=\frac{P_fV_f}{T_f},\\\space\\
V_f=V_i\frac{T_f}{T_i}."


(a) The final volume:


"V_f=V_i\frac{T_f}{T_i}=2·\frac{273+327}{273+27}=4\text{ m}^3."


(b) In this case:


"V_f=V_i\frac{T_f}{T_i}=2·\frac{273-123}{273+27}=1\text{ m}^3."